OpenStudy (anonymous):

word problem on Radical equation: Two added to the square root of the product of 4 and a number is equal to 10. A number is 6more than the principal square root of twice the number.

OpenStudy (anonymous):

2 + sqrt(4x) = 10 x = 16

OpenStudy (anonymous):

thanks and the second one?

OpenStudy (anonymous):

Not getting a resonable answer. the equation should be x-6 = sqrt(2x)

OpenStudy (anonymous):

It's fine guravS thanks agian. It'shard to do word problem unless you know a pattern for one

OpenStudy (anonymous):

\[x=6+\sqrt{2x}\] (x-6)^2=2x x^2-14x+36=0 \[x=7+\sqrt{13}, x=7-\sqrt{13}\]